# Observability of dynamical tides in merging eccentric neutron star binaries

### v1

@@ Abstract
While dynamical tides only become relevant during the last couple of orbits for circular inspirals, orbital eccentricity can increase their impact during earlier phases of the inspiral by exciting tidal oscillations at each close encounter. We investigate the effect of dynamical tides on the orbital evolution of eccentric neutron star binaries using post-Newtonian numerical simulations and constructing an analytic stochastic model. Our study reveals a strong dependence of dynamical tides on the pericenter distance, with the energy transferred to dynamical tides over that dissipated in gravitational waves (GWs) exceeding $\sim 1\%$ at separations $r_{\mathrm{p}}\lesssim 50$ km for large eccentricities. We demonstrate that the effect of dynamical tides on orbital evolution can manifest as a phase shift in the GW signal. We show that the signal-to-noise ratio of the GW phase shift can reach the detectability threshold of 8 with a single aLIGO detector at design densitivity for eccentric neutron star binaries at a distance of $40$ Mpc. This requires a pericenter distance of $r_{\mathrm{p0}}\lesssim 68$ km ($r_{\mathrm{p0}}\lesssim 76$ km) at binary formation with eccentricity close to 1 for a reasonable tidal deformability and f-mode frequency of 500 and $1.73$ kHz (700 and $1.61$ kHz), respectively. The observation of the phase shift will enable measuring the f-mode frequency of neutron stars independently from their tidal deformability, providing significant insights into neutron star seismology and the properties of the equation of state. We also explore the potential of distinguishing between equal-radius and twin-star binaries, which could provide an opportunity to reveal strong first-order phase transitions in the nuclear equation of state.
@@ I Introduction
Since the first direct observation of a binary black hole (BH) inspiral with gravitational waves (GWs) in September 2015 Abbott et al. (2016), the LIGO–Virgo–Kagra (LVK) Collaboration announced the detection of almost 100 GW signals produced by inspiraling compact object binaries Abbott et al. (2023), and the ongoing observing run already produced more than 117 additional significant candidate GW events111see https://gracedb.ligo.org/superevents/public/O4/ for the list of candidates. These events are dominated by binary BH mergers, however, some of them are attributed to mergers of binary neutron star (NS) and NS–BH systems. Particularly interesting was the GW event of a binary NS merger, detected in August 2017, GW170817 Abbott et al. (2017a). Together with the associated gamma-ray burst and kilonova signals they were the first example of multi-messenger astronomy with GWs Abbott et al. (2017b, c). This multi-messenger observation provided a rich laboratory for standard siren measurement of the cosmic expansion history Abbott et al. (2017d), the investigation of the equation of state (EoS) of strongly interacting matter through tidal deformability measurement Abbott et al. (2019), as well as the inference on the origin of r-process elements Kasen et al. (2017).
@@
While circular compact binary inspirals are expected to be more common, dynamical tides can be even more important in highly eccentric NS binary systems due to the broad spectrum of the acting tidal forces. During close encounters of the NSs, these tidal forces can excite the lowest-frequency f-mode oscillations of the NSs, which in turn influences the orbital evolution. Due to the broad spectrum of excitations, unlike in circular binaries, dynamical tides in eccentric binaries might become relevant well before the merger of the system, providing an opportunity to be observed Yang et al. (2018); Yang (2019); Vick and Lai (2019).
@@ II.1 The Newtonian solution
where $m_{\mathrm{tot}}=m_{1}+m_{2}$ and $\mu=m_{1}m_{2}/m_{\mathrm{tot}}$ are the total and reduced masses, $r$ and $\varphi$ are the relative distance and true anomaly, respectively, and $\omega_{0}$ is the fundamental f-mode frequency. Note that the static solution for the tidal deformation reads $Q_{ij}=-\lambda\mathcal{E}_{ij}$.
@@
For realistic NSs, $\gamma_{0}\ll\omega_{0}$ is satisfied. [^4] In our examples below, we typically assume $\lambda=\Lambda m_{\rm NS}^{5}$ where $400\leq\Lambda\leq 700$ for which $\gamma_{0}/\omega_{0}=\frac{1}{5}\Lambda(m_{\rm NS}\omega_{0})^{5}\sim 3\times 10^{-12}$ if $\omega_{0}=2\pi f_{0}$ and $1.61\,\mathrm{kHz}\leq f_{0}\leq 1.81\,\mathrm{kHz}$, see Sec. II.3.
One should in principle also account for the spins of the NSs, as they can strengthen dynamical tides Lai (1994); Ho and Lai (1999); Foucart et al. (2019); Ma et al. (2020); Steinhoff et al. (2021). If the NS is rotating in the retrograde direction with respect to the orbit, the rotation drags the oscillation mode frequencies to lower values in the inertial frame. This allows a resonance to occur at a lower orbital frequency. Studies investigating the effect of spins on dynamical tides have shown that low-frequency g-mode oscillations, which are weakly coupled to the external tides, can be enhanced if the NS rotates faster than $100$ Hz, while f-mode resonance can also significantly increase the phase shift if the spin frequency is higher than $300-500$ Hz, depending on the EoS Ho and Lai (1999); Lai and Wu (2006). However, most studies, as well as recent observations suggest that a high spin rate is unlikely for NS binaries at the time they enter the aLIGO band, due to spin-down allowed by the long timescale of the binary evolution. For example, both GW events associated with binary NS inspirals, GW170817 and GW190425, were found to be consistent with low spin configurations Abbott et al. (2019, 2020), while the fastest known spinning pulsar in a binary NS system spins with $44$ Hz Lyne et al. (2004). In less relativistic systems, such as e.g. white dwarf binaries, external torques acting on dynamical tides can spin up these objects in a way that they become tidally locked to the orbit Burkart et al. (2012, 2013, 2014). However, for binary NSs the spin-up rate is very low compared to the orbital frequency evolution and tidal locking fails to occur by many orders of magnitude at a wide range of orbital parameters Burkart et al. (2014); Lai (1994). Additionally, the spins of isolated field binaries are expected to be aligned with their orbital angular momentum (e.g. Ref. Farr et al. (2017)). Nevertheless, since high eccentricities and misaligned spins are both characteristic of dynamical formation channels, while the inspiral timescale of eccentric binaries are also much shorter than for circular ones, spin effects might still be important for these systems. For the time being, we neglect the effect caused by the rotation of NSs, however, this might be investigated in a follow-up study.
@@ III.1 Energy transmitted to tidal modes
However, we are interested in the energy exchange during a single passage. Then, the Fourier transform should be limited to only a single orbital period, while keeping the tidal field zero elsewhere. However, in many cases, especially close to the merger, where the tidal field at the apocenter is not negligible, the jump at the boundary results in an oscillatory spectrum, and therefore the energy exchange will also include this artificial oscillation. In fact, it is not trivial how to uniquely define the single-passage scenario without considering the whole evolution of the system. We find that a convenient way to define the energy transfer is to smoothly interpolate between the harmonics of the discrete Fourier spectrum. This way, we get a continuous and smooth spectrum, while this smooth interpolation also creates a natural smoothing at the edges of the tidal field in the time domain. This way we get the coefficient in Eq. (21) for all values of $\omega_{0}$ (for which $\omega_{0}=k\omega_{\rm orb}$ does not necessarily hold) by interpolating the discrete function:
@@
Since the tidal oscillations of the two NSs are independent of (not coupled to) each other, when taking into account the quadrupole deformation of both objects, the energies extracted from the orbit simply add up.
To understand the amount of energy transmitted to tidal modes during a single orbit, $\Delta E_{Q}$, it is useful to compare it to the amount of energy radiated away during a single orbit, $\Delta E_{\mathrm{GW}}=\dot{E}_{\mathrm{GW}}T_{\rm orb}$ (Eq. 18) where $T_{\mathrm{orb}}=2\pi/\omega_{\mathrm{orb}}$ is the orbital period. The solid black lines in Figs. 2 and 3 show the logarithm of $\Delta E_{Q}/\Delta E_{\mathrm{GW}}$ as a function of $\rho_{\mathrm{p}}$ and $e$ for a binary of two $1.4~{}\mathrm{M}_{\odot}$ NSs and one with such a NS and a $10~{}\mathrm{M}_{\odot}$ BH, respectively.
One can observe that there is an exponential-like dependency of $\Delta E_{Q}/\Delta E_{\mathrm{GW}}$ as a function of $\rho_{\mathrm{p}}$. This can be attributed to the fact that the Hansen coefficients have an exponential tail for large values of $k$, and increasing $\rho_{\mathrm{p}}$ means increasing $\omega_{0}/\omega_{\mathrm{orb}}$ as well. Looking at the numerical simulation results, the tidal interaction becomes visible above $\Delta E_{Q}/\Delta E_{\mathrm{GW}}\gtrsim 10^{-3}$. Another observation can be made by looking at the values of $T_{\mathrm{orb}}/\tau$, where $\tau=1/\gamma_{0}$ is the damping time. For most of the parameter space $T_{\mathrm{orb}}/\tau\ll 1$, which means that tidal oscillations do not have time to attenuate from one pericenter passage to the next. However, the oscillation may also be cancelled by the subsequent pulse-like excitation if it comes at the right phase. Due to the continuously changing orbital period, the phase at which the tidal pulse arrives is mostly random. Thus, tidal excitations follow a random walk and so will the energy extracted from the orbit. Similar tidal evolutions were found in Refs. Vick and Lai (2019); Samsing et al. (2017). The orange line in Fig. 2 shows the evolution when tidal effects are taken into account which deviate from simulations without tidal effects for small $\rho_{p}$. Note that this figure corresponds to NSs with tidal parameters $\Lambda=400$ and $f_{0}=1800$ Hz, and the excitation energy would be increased if $\Lambda$ was higher or $f_{0}$ lower.
@@
Figure 2: The logarithm of $\Delta E_{Q}/\Delta E_{\mathrm{GW}}$ as a function of orbital parameters $\rho_{\mathrm{p}}$ and $e$ for two $1.4$ $\mathrm{M}_{\odot}$ NSs (solid black lines). The dashed red lines show the orbit-averaged evolution of binaries with selected initial conditions $e_{0}\approx 1$ and $\rho_{\mathrm{p0}}=(10,15,20)$, driven by Eqs. 14-15. Black circles show the evolution of compact objects along the same line of evolution without any tidal interactions. These come from our numerical simulation, and each circle corresponds to an apocenter passage. The orange, light green and dark green dashed lines with circles correspond to compact binaries with the same initial parameters but with tidal interactions also taken into account. The tidal deformability and the f-mode frequency was chosen as $\Lambda=400$ and $f_{0}=\omega_{0}/2\pi=1800$ Hz. The blue line represents parameters at the last stable orbit, while the vertical orange line corresponds to direct collision of the two NSs, where we used the very conservative estimate $R_{\mathrm{NS}}(1.4~{}\mathrm{M}_{\odot})=14$ km. Dashed black lines show different values of $T_{\mathrm{orb}}/\tau$. For each selected orbital evolution we display the sky position and orientation averaged detection range with a single aLIGO detector, $d_{\rm L,det}$, calculated from the orbit-averaged solution O’Leary et al. (2009).
Figure 3: Same as Fig. 2 but for a binary consisting of a NS with $m_{1}=1.4~{}\mathrm{M}_{\odot}$ and a BH with $m_{2}=10~{}\mathrm{M}_{\odot}$. The three selected orbital evolutions here have $e_{0}\approx 1$ and $\rho_{\mathrm{p0}}=(5,6,8)$. Black circles corresponding to an evolution without any tidal interaction were omitted, since they would overlap with the colored circles. The tidal parameters of the NS are $\Lambda=400$ and $f_{0}=1800$ Hz.
@@ III.2 Phase shift caused by tidal interactions
Now we also need to take into account, that the amplitude of tidal excitations follows a random walk due to the random phase at which the short tidal pulses arrive. The $j^{\mathrm{th}}$ pulse can be characterized by the complex number $\Delta q_{j}=A_{j}\mathrm{exp}(i\varphi_{j})$, where $\varphi_{i}$ is a random phase in the $[0,2\pi]$ interval. Thus, the expectation value of the complex amplitude after the $j^{\mathrm{th}}$ pulse is
@@ III.3 Comparison to numerical simulations
Fig. 6 shows the results obtained from this procedure for binary NSs with two $1.4~{}\mathrm{M}_{\odot}$ components and three different configurations of initial semi-major axis and tidal deformability. The initial eccentricity is $e_{0}=0.9$ for all cases. The solid lines represent the semi-analytical results obtained from Eq. (30), utilizing the numerically calculated orbital evolution without dynamical tides, while the dashed lines represent the numerical results obtained from the procedure described in the previous paragraphs. For each combination of binary parameters, two results from numerical simulations with slightly different parameters are shown. Due to the random walk of phase shifts between consecutive orbits, the difference between the two evolutions can become considerable, however, they both follow the trend given by the semi-analytic formula. Even though the magnitude of the phase shifts for the three different configurations differs by $2-3$ orders of magnitude at low orbital frequencies, their difference shrinks to a single order of magnitude by the merger. The phase shifts approach or even exceed 1 radian by the time the binaries merge for all three configurations.
@@
As we mentioned in the beginning of this subsection, a correct treatment of the post-Newtonian interactions would require taking into account additional corrections to equations of motion used in the numerical simulations and the analytical formulas. For a simple test, we estimate the change in the phase shift evolution caused solely by the 1PN and 2PN terms added to the orbital evolution equations (see Eqs. 53 and 54. Fig. 8 shows the results for the phase shift as a function of the orbital frequency when the 1PN and 2PN terms are included in the orbital equations of motion. The numerical evaluation of the phase shift is described in Appendix B. $f_{\mathrm{orb}}$ was calculated using the PN corrections described in Ref. Blanchet (2014). There is a good overall agreement between the analytic and numerical results, one major difference being that the phase shifts for $a_{0}=800\,$km and $\Lambda=700$ are systematically increased compared to the analytic results. Since the relative scatter of phase shifts is also reduced in this case, we think that it might be caused by an additional, non-stochastic effect, although the exact reason remains unclear. It should also be noted that with the inclusion of the 1PN and 2PN corrections, the orbital frequency at merger is reduced by about a factor of two, resulting in a lower phase shift at merger, while the total number of orbits is also reduced.

### v2

@@ Abstract
While dynamical tides only become relevant during the last couple of orbits for circular inspirals, orbital eccentricity can increase their impact during earlier phases of the inspiral by exciting tidal oscillations at each close encounter. We investigate the effect of dynamical tides on the orbital evolution of eccentric neutron star binaries using post-Newtonian numerical simulations and construct an analytic stochastic model that reproduces the numerical results. Our study reveals a strong dependence of dynamical tides on the pericenter distance, with the fractional energy transferred to dynamical tides over that dissipated in gravitational waves (GWs) exceeding $\sim 1\%$ at separations $r_{\mathrm{p}}\lesssim 50$ km for large eccentricities. We demonstrate that the effect of dynamical tides on orbital evolution can manifest as a phase shift in the GW signal. We show that the signal-to-noise ratio of the GW phase shift can reach the detectability threshold of 8 with a single aLIGO detector at design densitivity for eccentric neutron star binaries at a distance of $40$ Mpc. This requires a pericenter distance of $r_{\mathrm{p0}}\lesssim 68$ km ($r_{\mathrm{p0}}\lesssim 76$ km) at binary formation with eccentricity close to 1 for a reasonable tidal deformability and f-mode frequency of 500 and $1.73$ kHz (700 and $1.61$ kHz), respectively. The observation of the phase shift will enable measuring the f-mode frequency of neutron stars independently from their tidal deformability, providing significant insights into neutron star seismology and the properties of the equation of state. We also explore the potential of distinguishing between equal-radius and twin-star binaries, which could provide an opportunity to reveal strong first-order phase transitions in the nuclear equation of state.
@@ I Introduction
Since the first direct observation of a binary black hole (BH) inspiral with gravitational waves (GWs) in September 2015 Abbott et al. (2016), the LIGO–Virgo–Kagra (LVK) Collaboration Martynov et al. (2016); Acernese et al. (2019); Akutsu et al. (2021) announced the detection of almost 100 GW signals produced by inspiraling compact object binaries Abbott et al. (2023), and the ongoing observing run already produced more than 117 additional significant candidate GW events111see https://gracedb.ligo.org/superevents/public/O4/ for the list of candidates. These events are dominated by binary BH mergers, however, some of them are attributed to mergers of binary neutron star (NS) and NS–BH systems. Particularly interesting was the GW event of a binary NS merger, detected in August 2017, GW170817 Abbott et al. (2017a). Together with the associated gamma-ray burst and kilonova signals they were the first example of multi-messenger astronomy with GWs Abbott et al. (2017b, c). This multi-messenger observation provided a rich laboratory for standard siren measurement of the cosmic expansion history Abbott et al. (2017d), the investigation of the equation of state (EoS) of strongly interacting matter through tidal deformability measurement Abbott et al. (2019), as well as the inference on the origin of r-process elements Kasen et al. (2017).
@@
While circular compact binary inspirals are expected to be more common, dynamical tides can be even more important in highly eccentric NS binary systems due to the broad spectrum of the acting tidal forces. During close encounters of the NSs, these tidal forces can excite the lowest-frequency f-mode oscillations of the NSs, which in turn influences the orbital evolution. Due to the broad spectrum of excitations, unlike in circular binaries, dynamical tides in eccentric binaries might become relevant well before the merger of the system, providing an opportunity to be observed Yang et al. (2018); Yang (2019); Vick and Lai (2019); Gamba and Bernuzzi (2023); Chaurasia et al. (2018).
@@ II.1 The Newtonian solution
where $m_{\mathrm{tot}}=m_{1}+m_{2}$ and $\mu=m_{1}m_{2}/m_{\mathrm{tot}}$ are the total and reduced masses, and $\omega_{0}$ is the fundamental f-mode frequency. The Lagrangian is expressed in the Newtonian inertial frame, with $r=|\bm{r}|\equiv|\bm{x}_{1}-\bm{x}_{2}|$ being the relative distance of the two objects and $\varphi$ is the true longitude, i.e. the angle between $\bm{e}_{x}$ and $\bm{r}$. Note that the static solution for the tidal deformation reads $Q_{ij}=-\lambda\mathcal{E}_{ij}$.
@@
For realistic NSs, $\gamma_{0}\ll\omega_{0}$ is satisfied. [^4] In our examples below, we typically assume $\lambda=\Lambda m_{\rm NS}^{5}$ where $400\leq\Lambda\leq 700$ for which $\gamma_{0}/\omega_{0}=\frac{1}{5}\Lambda(m_{\rm NS}\omega_{0})^{5}\sim 3\times 10^{-4}$ if $\omega_{0}=2\pi f_{0}$ and $1.61\,\mathrm{kHz}\leq f_{0}\leq 1.81\,\mathrm{kHz}$, see Sec. II.3.
One should in principle also account for the spins of the NSs, as they can strengthen dynamical tides Lai (1994); Ho and Lai (1999); Foucart et al. (2019); Ma et al. (2020); Steinhoff et al. (2021); Yu et al. (2024). If the NS is rotating in the retrograde direction with respect to the orbit, the rotation drags the oscillation mode frequencies to lower values in the inertial frame. This allows a resonance to occur at a lower orbital frequency. Studies investigating the effect of spins on dynamical tides have shown that low-frequency g-mode oscillations, which are weakly coupled to the external tides, can be enhanced if the NS rotates faster than $100$ Hz, while f-mode resonance can also significantly increase the phase shift if the spin frequency is higher than $300-500$ Hz, depending on the EoS Ho and Lai (1999); Lai and Wu (2006). However, most studies, as well as recent observations suggest that a high spin rate is unlikely for NS binaries at the time they enter the aLIGO band, due to spin-down allowed by the long timescale of the binary evolution. For example, both GW events associated with binary NS inspirals, GW170817 and GW190425, were found to be consistent with low spin configurations Abbott et al. (2019, 2020), while the fastest known spinning pulsar in a binary NS system spins with $44$ Hz Lyne et al. (2004). In less relativistic systems, such as e.g. white dwarf binaries, external torques acting on dynamical tides can spin up these objects in a way that they become tidally locked to the orbit Burkart et al. (2012, 2013, 2014). However, for binary NSs the spin-up rate is very low compared to the orbital frequency evolution and tidal locking fails to occur by many orders of magnitude at a wide range of orbital parameters Burkart et al. (2014); Lai (1994). Additionally, the spins of isolated field binaries are expected to be aligned with their orbital angular momentum (e.g. Ref. Farr et al. (2017)). Nevertheless, since high eccentricities and misaligned spins are both characteristic of dynamical formation channels, while the inspiral timescale of eccentric binaries are also much shorter than for circular ones, spin effects might still be important for these systems. For the time being, we neglect the effect caused by the rotation of NSs, however, this might be investigated in a follow-up study.
@@ III.1 Energy transmitted to tidal modes
However, we are interested in the energy exchange during a single passage. Then, the Fourier transform should be limited to only a single orbital period, while keeping the tidal field zero elsewhere. However, in many cases, especially close to the merger, where the tidal field at the apocenter is not negligible, the jump at the boundary results in an oscillatory spectrum, and therefore the energy exchange will also include this artificial oscillation. In fact, it is not trivial how to uniquely define the single-passage scenario in the moderate and low-eccentricity case, where tidal forces in the apocenter are non-negligible, without considering the entire evolution of the system. We find that a convenient way to define the energy transfer is to smoothly interpolate between the harmonics of the discrete Fourier spectrum. This way, we get a continuous and smooth spectrum, while this smooth interpolation also creates a natural smoothing at the edges of the tidal field in the time domain. We can then obtain the coefficient in Eq. (21) for all values of $\omega_{0}$ (for which $\omega_{0}=k\omega_{\rm orb}$ does not necessarily hold) by interpolating the discrete function:
@@
Since the tidal oscillations of the two NSs are independent of (not coupled to) each other, when taking into account the quadrupole deformation of both objects, the energies extracted from the orbit simply add up. Comparing to our numerical simulations, we find that this formula remains accurate (with a relative error of $\lesssim 10\%$) even for eccentricities as low as $e\approx 0.2$.
To understand the amount of energy transmitted to tidal modes during a single orbit, $\Delta E_{Q}$, it is useful to compare it to the amount of energy radiated away during a single orbit, $\Delta E_{\mathrm{GW}}=\dot{E}_{\mathrm{GW}}T_{\rm orb}$ (Eq. 18) where $T_{\mathrm{orb}}=2\pi/\omega_{\mathrm{orb}}$ is the orbital period. The color maps in Figs. 2 and 3 show the logarithm of $\Delta E_{Q}/\Delta E_{\mathrm{GW}}$ as a function of $\rho_{\mathrm{p}}$ and $e$ for a binary of two $1.4~{}\mathrm{M}_{\odot}$ NSs and one with such a NS and a $10~{}\mathrm{M}_{\odot}$ BH, respectively.
One can observe that there is an exponential-like dependency of $\Delta E_{Q}/\Delta E_{\mathrm{GW}}$ as a function of $\rho_{\mathrm{p}}$. This can be attributed to the fact that the Hansen coefficients have an exponential tail for large values of $k$, and increasing $\rho_{\mathrm{p}}$ means increasing $\omega_{0}/\omega_{\mathrm{orb}}$ as well. Looking at the numerical simulation results, the tidal interaction becomes visible above $\Delta E_{Q}/\Delta E_{\mathrm{GW}}\gtrsim 10^{-3}$. Another observation can be made by looking at the values of $T_{\mathrm{orb}}/\tau$, where $\tau=1/\gamma_{0}$ is the damping time. For most of the parameter space $T_{\mathrm{orb}}/\tau\ll 1$, which means that tidal oscillations do not have time to attenuate from one pericenter passage to the next. However, the oscillation may also be cancelled by the subsequent pulse-like excitation if it comes at the appropriate phase. Due to the continuously changing orbital period, the phase at which the tidal pulse arrives is seemingly random. Thus, tidal excitations will appear to follow a random walk and so will the energy extracted from the orbit. Similar tidal evolutions were found in Refs. Vick and Lai (2019); Samsing et al. (2017). The red lines with markers in Fig. 2 show orbital evolutions when tidal effects are taken into account which deviate from simulations without tidal effects (dashed black lines) for small $\rho_{p}$. Note that this figure corresponds to NSs with tidal parameters $\Lambda=400$ and $f_{0}=1800$ Hz, and the excitation energy would be increased if $\Lambda$ was higher or $f_{0}$ lower.
@@
Figure 2: The logarithm of $\Delta E_{Q}/\Delta E_{\mathrm{GW}}$ as a function of orbital parameters $\rho_{\mathrm{p}}$ and $e$ for two $1.4$ $\mathrm{M}_{\odot}$ NSs. The dashed black lines show the orbit-averaged evolution of binaries with selected initial conditions $e_{0}\approx 1$ and $\rho_{\mathrm{p0}}=(10,15,20)$, driven by Eqs. 14-15. The dashed red lines with different markers correspond to compact binaries with the same initial parameters but with tidal interactions also taken into account. These come from our numerical simulation, and each marker corresponds to an apocenter passage. The tidal deformability and the f-mode frequency was chosen as $\Lambda=400$ and $f_{0}=\omega_{0}/2\pi=1800$ Hz. The blue line represents parameters at the last stable orbit, while the vertical yellow line corresponds to direct collision of the two NSs, where we used the very conservative estimate $R_{\mathrm{NS}}(1.4~{}\mathrm{M}_{\odot})=14$ km. Dashed green lines show different values of $T_{\mathrm{orb}}/\tau$. For each selected orbital evolution we display the sky position and orientation averaged detection range with a single aLIGO detector, $d_{\rm L,det}$, calculated from the orbit-averaged solution O’Leary et al. (2009).
Figure 3: Same as Fig. 2 but for a binary consisting of a NS with $m_{1}=1.4~{}\mathrm{M}_{\odot}$ and a BH with $m_{2}=10~{}\mathrm{M}_{\odot}$. The three selected orbital evolutions here have $e_{0}\approx 1$ and $\rho_{\mathrm{p0}}=(5,6,8)$. The tidal parameters of the NS are $\Lambda=400$ and $f_{0}=1800$ Hz.
@@ III.2 Phase shift caused by tidal interactions
Now we also need to take into account that while the evolution of the tides is fundamentally deterministic, due to the seemingly random phase at which the short tidal pulses arrive, the amplitude of tidal excitations will appear to vary stochastically. The $j^{\mathrm{th}}$ pulse can be characterized by the complex number $\Delta q_{j}=A_{j}\mathrm{exp}(i\varphi_{j})$, where $\varphi_{i}$ is a random phase in the $[0,2\pi]$ interval. Thus, the expectation value of the complex amplitude after the $j^{\mathrm{th}}$ pulse is
@@ III.3 Comparison to numerical simulations
Fig. 6 shows the results obtained from this procedure for binary NSs with two $1.4~{}\mathrm{M}_{\odot}$ components and three different configurations of initial semi-major axis and tidal deformability. The initial eccentricity is $e_{0}=0.9$ for all cases. The solid lines represent the semi-analytical results obtained from Eq. (30), utilizing the numerically calculated orbital evolution without dynamical tides, while the dashed lines represent the numerical results obtained from the procedure described in the previous paragraphs. For each combination of binary parameters, two results from numerical simulations with slightly different parameters are shown. Due to the stochastic motion of phase shifts between consecutive orbits, the difference between the two evolutions can become considerable, however, they both follow the trend given by the semi-analytic formula. Even though the magnitude of the phase shifts for the three different configurations differs by $2-3$ orders of magnitude at low orbital frequencies, their difference shrinks to a single order of magnitude by the merger. The phase shifts approach or even exceed 1 radian by the time the binaries merge for all three configurations.
@@
As we mentioned in the beginning of this subsection, a correct treatment of the post-Newtonian interactions would require taking into account additional corrections to equations of motion used in the numerical simulations and the analytical formulas. For a simple test, we estimate the change in the phase shift evolution caused solely by the 1PN and 2PN terms added to the orbital evolution equations (see Eqs. 55 and 56. Fig. 8 shows the results for the phase shift as a function of the orbital frequency when the 1PN and 2PN terms are included in the orbital equations of motion. The numerical evaluation of the phase shift is described in Appendix B. $f_{\mathrm{orb}}$ was calculated using the PN corrections described in Ref. Blanchet (2014). There is a good overall agreement between the analytic and numerical results, one major difference being that the phase shifts for $a_{0}=800\,$km and $\Lambda=700$ are systematically increased compared to the analytic results. Since the relative scatter of phase shifts is also reduced in this case, we think that it might be caused by an additional, non-stochastic effect, although the exact reason remains unclear. It should also be noted that with the inclusion of the 1PN and 2PN corrections, the orbital frequency at merger is reduced by about a factor of two, resulting in a lower phase shift at merger, while the total number of orbits is also reduced.